The formula unit mass of X is 342 u. Among the following, X is :

Grade:12th pass

3 Answers

Pawan Prajapati
2 years ago
Our expert is working on this class X maths answer. We will update the answer very soon.
Yash Patari
2 years ago

Dear student

Formula unit mass of Al2
(SO4
)33=
(2× Atomic mass of Al)+(3×Mass of S)+(12× Mass of O
)

=2×27 u+3×32 u+12×16 u
=342 u

Thanks
Yash Patari
2 years ago
(iiFormula unit mass of any compound is the sum of atomic masses of all atoms in the formula of that compound.  Formula unit mass of Al2(SO4)3 
Atomic mass of Al = 27 × 2 = 54 u
  Atomic mass of S = 32 × 3 = 96 u 
Atomic mass of O = 16 × 12 = 192 u
  Formula unit mass = 54 + 96 + 192 = 342 u.  Hence, Formula unit mass of Al2(SO4)3 is 342.